OpenStudy (nightmarenight):

Help? Will fan and give medal for right answer!! If f(x) varies directly with x^2, and f(x) = 75 when x = 5, find the value of f(4)

OpenStudy (nikato):

So f(x) is like y. So ur equation can be written as y=x^2 (k). Where k is a constant which u have to find. So find k when y=75 and x=5 75=5^2 (k) Can you solve for k?

OpenStudy (nightmarenight):

25? @nikato

OpenStudy (nikato):

No. 75=25k. Divide both side by 25. So what's k?

OpenStudy (nightmarenight):

3 @nikato

OpenStudy (nightmarenight):

3=1 @nikato

OpenStudy (nikato):

ok. so k=3 now ur equation is y=x^2(3) and now ur solving for f(x) which means that x=4 plug 4 into x and solve for y y=4^2(3)

OpenStudy (nightmarenight):

24 @nikato

OpenStudy (nightmarenight):

@nikato

OpenStudy (nikato):

now what is 4^2? then multiply that by 3

OpenStudy (nightmarenight):

16 x 3= 48 lol sorry i was doing 4x2 @nikato

OpenStudy (nikato):

no problem. 48 is correct